# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 4.2.1 Correctness and Complexity
If no agent gets consumed by the byzantine black hole, while performing an iteration of Algorithm 2, then our algorithm ensures that in that iteration, either the ring $R$ is explored or there is one agent that knows exactly the location of the black hole.
Proof 4.16.
Suppose till the completion of $i$-th iteration $(i>0)$ no agent is consumed by the black hole, then in any iteration $0<j\leq i$, the agent $a_{1}$ follows the sequence $SEQ1$, whereas $a_{2}$ follows the sequence $SEQ2$, as instructed starting from the initial $home$ node. Considering this scenario we have two cases:
  • No agents and no pebbles are destroyed: In this case, while executing the sequence $SEQ1$, the agent after changing its state to Leader from Initial, visits a new node after every 4 rounds in clockwise direction. Since, no agent has been consumed in this iteration as well, so $a_{1}$ un-obstructively visits each node in a clockwise direction, and reaches $home$ to finally end this sequence $SEQ1$, while again changing its state to Initial. This shows that, at least one agent visits each node of $R$ in an iteration completing the exploration of $R$.
  • No agent gets destroyed but some pebbles are consumed by the black hole: Note that, this situation can only occur when the black hole consumes the pebble carried and released by $a_{2}$, while $a_{2}$ is not at the node containing the pebble, executing the sequence $SEQ2$. It is because, in an iteration, only $a_{1}$ and $a_{2}$ moves away from $home$ while executing their respective sequences $SEQ1$ and $SEQ2$. And the pebble carried by $a_{1}$ ( i.e., the agent with smallest ID) cannot be destroyed without destroying $a_{1}$, as $a_{1}$ always carries the pebble at each node while executing its respective sequence. So, in this situation, while $a_{2}$ ( i.e., the agent with second lowest ID) in state Follower-Collect moves back to the adjacent node in counter-clockwise direction, in order to collect the already released pebble, absence of which triggers $a_{2}$ to conclude the current node to be the black hole, and which leads the agent to immediately leave the black hole node. In this case, the consumption of a pebble leads to the black hole detection.
@@
If no agent gets consumed by the byzantine black hole while performing an iteration of Algorithm 2 but some pebbles are consumed by the black hole, then by Lemma 4.15, one agent (more precisely, the agent with second lowest ID, i.e., $a_{2}$) detects the exact location of the black hole. Further, this agent continues to perpetually explore the ring $R$ by avoiding the black hole node.
@@
    The exact location of the black hole is detected by at least one agent.
@@
    All alive agents become colocated and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which is the black hole.
Proof 4.18.
@@
Note that the agent which has entered the black hole is either $a_{1}$ or $a_{2}$, based on which we have the following conditions:
@@
  • $a_{1}$ falls in to the black hole: This situation can only occur when $a_{1}$ following sequence $SEQ1$ is in state Leader, visits the black hole node along with its pebble. Now, based on the adversarial choice of byzantine black hole, the pebble may or may not be destroyed by the black hole. Irrespective of which, $a_{2}$ within 3 additional rounds reaches the black hole node while in state Follower-Find, and absence of $a_{1}$ triggers the agent $a_{2}$ to determine the current node as black hole, and immediately leave the current node.
@@
  • $a_{2}$ falls in to the black hole: This can only happen when $a_{2}$ is not with $a_{1}$ at the black hole node. Otherwise, both agents will be destroyed, contradicting our claim. Note that, $a_{2}$ can either be with the pebble or alone, while it is consumed by the black hole. Also note that, since $a_{2}$ is executing the sequence $SEQ2$, hence the distance between $a_{1}$ and $a_{2}$ can be at most 2. So during an iteration, if $a_{2}$ is destroyed by a black hole, then the black hole must be any one of the two consecutive counter-clockwise nodes, from the current position of $a_{1}$. If the current node is not $home$, in this case, $a_{1}$ gets to know about this fact, only when it finds the absence of $a_{2}$ even after waiting for 3 rounds, at the current node. Now, according to our algorithm, it leaves the pebble accompanied by it, and moves from the current node in clockwise direction to $home$ with state Report-Leader. After, reaching $home$, it waits until the end of this iteration, i.e., until $4n$ rounds from the start of this iteration and moves to state Find-Pebble. During the $14n+$ 1-th round from the start of the current iteration, $a_{3}$ finds that it is only with $a_{1}$, whereas only one pebble exits, this leads the agent to change its state to Find-Pebble. In this state, both the agents move counter-clockwise from $home$ until they encounter a pebble. Note that, this pebble is the one left by $a_{1}$ after determining the fact that $a_{2}$ has entered the black hole. So, now both these agents $a_{1}$ and $a_{3}$, declare the node with pebble as $home$, whereas also denote the adjacent counter-clockwise node from the current node as $S_{lft}$, whereas the other node adjacent to $S_{lft}$ in counter-clockwise direction as $S_{rgt}$. On the other hand, if the current node of $a_{1}$ is $home$, while $a_{2}$ is destroyed by the black hole, then $a_{2}$ can never reach $home$ along with the pebble it was carrying by following the sequence $SEQ2$. So, at the end of this iteration, both $a_{1}$ and $a_{3}$ finds that there are two agents (i.e., $a_{1}$ and $a_{3}$) and two pebbles, respectively at $home$. This leads both of them to conclude that, the adjacent counter-clockwise node to be $S_{rgt}$, whereas the adjacent counter-clockwise node of $S_{rgt}$ as $S_{lft}$.
So, in each case we showed that either of our two conditions hold.
@@
Following from Lemma 4.17, if exactly one agent gets destroyed by the byzantine black hole and the exact location of the black hole is detected by at least one agent, then our algorithm ensures that the ring is perpetually explored by at least one alive agent while avoiding the black hole node. On the contrary, all alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, among which one is a black hole, then until and unless another agent falls in to the black hole, our algorithm ensures that the ring is perpetually explored, by the remaining alive agents. Note that, by Lemma 4.20 while any one among them gets destroyed, then the other agent detects the exact black hole node, and then further perpetually explores the ring $R$, avoiding the black hole node.
@@
Algorithm PerpExplore-Coloc-Pbl ensures that if two among three agents gets consumed by the byzantine black hole, then the remaining agent knows the exact location of the black hole within some additional finite rounds without being destroyed.
Proof 4.21.
@@
Let us suppose, by contradiction, there exists a round $r$ ($>0$), within which two agents gets destroyed by the black hole, whereas the alive agent is unable to detect the exact location. Now we have the following cases:
@@
  • Both agents gets consumed by the black hole in the same iteration: Note that in this case, the consumed agents are $a_{1}$ and $a_{2}$, respectively, which while executing the sequences $SEQ1$ and $SEQ2$ enters the black hole. Also note that, $a_{1}$ and $a_{2}$ can only be together at a node, when $a_{2}$ has left the pebble it is carrying in the counter-clockwise node. So, now $a_{3}$ after the end of this iteration, finds that no agent among $a_{1}$ and $a_{3}$ has reached $home$, which triggers $a_{3}$ to change to state Find-BH. In this state, it moves clockwise, until it finds a pebble (this pebble is the one left by $a_{2}$ in the counter-clockwise node of the black hole). Whenever a pebble is found, $a_{3}$ concludes that the next node is the black hole, which contradicts our claim.
@@
  • There exists an iteration where exactly one agent is destroyed and the other one is destroyed during round $r$: Now in this case let $r^{\prime}<r$ be a round when exactly an agent is destroyed during an iteration, say $i>0$. Then by Lemma 4.17, within finitely many additional rounds $r_{0}$, where $r_{0}+r^{\prime}<r$, both the alive agents becomes colocated and agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which must be the black hole. Observe that according to our algorithm, the node which is clockwise adjacent to $S_{rgt}$ is determined to be the $home$ (if not already so). Note that the other case according to Lemma  4.17 where after exactly one agent is destroyed by the black hole during an iteration at a round $r^{\prime}<r$, there is an agent that knows the exact location of the black hole within a finite additional rounds and the agent starts exploring the ring avoiding the black hole and so even after round $r$ it stays alive knowing the exact location of the black hole contradicting our assumption. Now, when both the agents are colocated during a round say $r_{0}+r^{\prime}<r$ and knows that the black hole must be any of the two consecutive nodes $S_{lft}$ or, $S_{rgt}$. In this case, they move in to state Detection. After which one agent moves clockwise until it reaches $S_{lft}$ and the other one moves counter clockwise until it reaches $S_{right}$ and moves back to $home$. The agent that visits $S_{rgt}$ waits for the other agent at $home$ for $62n-$ 6 rounds which is sufficient for the other agent to return back at $home$. Now when during round $r$, another agent among these two gets destroyed, within at most $(2n-4)$ rounds they fails to meet at $home$. Which trigers the remaining alive agent to determine exactly which one among $S_{lft}$ or, $S_{rgt}$ is the black hole. This also contradicts the claim we made earlier. Thus, The algorithm PerpExplore-Coloc-Pbl ensures that, if two agent among three agents gets consumed by the byzantine black hole, the remaining agent must know the exact location of the black hole in some finite additional rounds without getting destroyed.
Corollary 4.22.
Number of alive agents must be at least 2, when the exact position of the byzantine black hole is not known to them.
From Lemma 4.20 and algorithm PerpExplore-Coloc-Pbl we have the following theorem
Theorem 4.23.
A set of 3 synchronous colocated agents equipped with one pebble each are necessary and sufficient to solve PerpExploration-BBH on a ring $R$.
Remark 4.24 (Algorithm on whiteboard model of communication).
The PerpExplore-Coloc-Pbl algorithm (i.e., Algorithm 2 ) also works for colocated agents with whiteboard model of communication. The agents can easily simulate a pebble on a node by a bit marked on that nodes memory. Also, carrying a pebble from one node to another can be simulated by removing the marking from the current node’s memory and move to a new node where the agent writes a new bit. Thus 3 colocated agents are also sufficient to solve Perpexplore-BBH when the nodes are equipped with whiteboards.
5 Scattered Agents
5.1 Communication: Pebble
@@
If 4 agents are scattered in more than one nodes initially then there are 3 cases. Firstly, all four agents are at different nodes. Secondly, Four agents are scattered in three different nodes nodes initially. For this case there exists exactly one node with two agents and remaining two nodes with exactly one agents each. In the third case, four agents are scattered in two nodes initially. In this case either each of the two nodes contains exactly two agents or, one node contains three agents and the other one has exactly one agent. We here describe and present the pseudocode considering the first case only. The case with 3 agents in one node can be dealt by instructing the colocated agents to execute PerpExplore-Coloc-Pbl when there are three agents on the current node (i.e., from the beginning). The algorithm we discuss here can also be used for the remaining cases with slight modifications. These modifications are described in the Remark 5.7
@@
The main idea of the PerpExplore-Scat-Pbl algorithm is that 4 agents perpetually explore the ring, when no agents are consumed by the black hole. Our algorithm ensures that during this exploration at most one agent can be destroyed. In this scenario, the remaining agents within further finite time, gather at a single node (which is the starting node of either of these 4 agents, hence it is safe), and after which they together start executing the algorithm PerpExplore-Coloc-Pbl.
@@
At the beginning of the Algorithm 3 all agents are in state Initial1. In this state, the agents declare their current node as $home$, moreover they initialize the variables $size=0$ and $s=0$, where $size$ variable is used by the agents in order to store the length of the path between it’s own $home$ and the nearest $home$ of another agent in the clockwise direction. Further the variable $s$ indicates the state transition. If the agent is at state Forward and $s=$ 0 then the agent moved from state Initial1 to its current state (i.e., state Forward), otherwise if the agent is at state Forward and $s=1$ then this indicates it changes to state Forward from state Wait2. An agent $a_{1}$ in state Forward, initializes $W_{time}=0$ (which stores the waiting time) and checks whether $s=0$ or $s=1$. Note that, if $s=0$ and the current node is $home$, then the $a_{1}$ is first instructed to leave the pebble at current node and then move clockwise direction, after increasing $size$ by 1. The next node where $a_{1}$ has reached can either be $home$ of another agent, say $a_{2}$, or it is not $home$ for any other agents. For the first case $a_{1}$ would find a pebble on its $Current-Node$ which was left by $a_{2}$. In this case $a_{1}$ updates $s$ to 1 and moves into state Wait1. On the other hand, if the $Current-Node$ is not $home$ of any other agents then $a_{1}$ finds no pebble at the $Current-Node$ and thus moves clockwise after increasing the variable $size$ by 1. Note that, when $s=0$ and an agent is in state Forward it always finds another pebble on some node, if it is not destroyed (this pebble must be left by another agent which is simultaneously executing Algorithm 3 and currently in state Forward). This node must be the clockwise nearest $home$ (i.e., $home$ of another agent which is clockwise nearest to own $home$). Now since during each move at state Forward, an agent increases the variable $size$ by 1, the $size$ variable must store the length between an agents own $home$ and the clockwise nearest $home$ when the agent finds another pebble. Beyond this point (i.e., when $s=1$) the agents will use the $size$ variable as reference to calculate how much it should move, as after this, as there is a possibility that the agent may not find a pebble at the clockwise nearest $home$.
@@
In state Wait2, the agent first checks whether $W_{time}<2n-size$, if so then it further checks whether #$\#Agent$ at current node is exactly 1. If these two conditions are satisfied then the agent just increments $W_{time}$ by 1. On the other hand, if it finds that $W_{time}<2n-size$ but #$\#Agent$ at the current node is 2, then it changes its state to Gather2 and updates $W_{time}$ to 0. Otherwise, if it finds $W_{time}<2n-size$ but #$\#Agent$ at the current node is 3 then it updates $W_{time}$ to 0, whereas changes its state to Coloc. Lastly, whenever it finds $W_{time}=2n-size$, it changes its state to Forward.
Note that if no agent gets destroyed by the black hole, then our algorithm guarantees that an agent can neither changes its state to Gather1 or in state Gather2. It is because, an agent, say $a_{1}$ can only find more than one pebble at its $home$ node, only when another agent say $a_{2}$ (exploring the counter-clockwise nodes of $home$ of $a_{1}$), which was supposed to fetch the pebble left by $a_{1}$ at its $home$, is destroyed. In this case, whenever $a_{1}$ returns while in state fetch with another pebble, it finds more than one pebble. So the first agent, which first finds this anomaly changes its state to Gather1 from the state Fetch. Now this agent must initiate gathering with the remaining two agents (as it is the one which has first detected the anomaly of more than one pebble at its $home$). Note that, if while $a_{1}$ starts its gathering phase, moving in a clockwise direction, there is a possibility that other agents are still moving either in clockwise or counter-clockwise direction. So, to address this challenge, the algorithm instructs $a_{1}$ (i.e., the agent which is in state Gather1) to wait at the current node (i.e., its $home$) for precisely $n-size-1$ rounds, and then start moving clockwise direction. This waiting time of $a_{1}$ guarantees all the other alive agents to reach their respective $home$ while in state Fetch. Now, in order to avoid further movement of remaining agents (i.e., again change to state Forward, since these agents are yet to detect any anomaly, which makes gathering a bit challenging), our algorithm instructs the remaining agents to wait at the current node for a certain number of rounds (i.e., more precisely, $12n-size-$ 1). This waiting period is sufficient, for the agent $a_{1}$ to meet with the other alive agents while in state Gather1. Note that, if no agent detects any anomaly, then they directly change their state to Wait2 from the state Fetch. If no such agent exists which encounters any anomaly and changes its state to Gather1 or Gather2 then this waiting time of $12n-size-$ 1 rounds, also ensures that all of them must move to state Forward at the exact same round (i.e, ($23n+$ 2)-th round from the previous time it moved to state Forward) [^1] At $(n+1)$-th round an agent moves to state Fetch after that it moves for $size$ rounds to reach its $home$ in counter-clockwise direction and changes to state Wait2. So at round $n+3+size$ an agent changes to state Wait2. After that waits for further $12n-size-$ 1 rounds and changes to state forward, taking a total of $23n+$ 2 rounds . Now we have the following observation {observation} Let $t$ be a round when all alive agents moved into state Forward. If no agent detects any anomaly (i.e., no agent finds more than one pebble at home) then, at round $t+3n+2$ all the agents together move into state Forward again.
An agent in state Gather1, is the first agent to detect the anomaly, hence it is first instructed to wait at the current node, i.e., its $home$ for the first $n-size-1$ rounds, and then start moving in a clockwise direction, while accompanying all the pebbles at the current node  [^2] Note that we have considered that in this case an agent, while in state Gather1 or Gather2 can carry more than one pebble, which can be easily restricted to one agent can carry only one pebble, but in that case the agent needs to return and carry one pebble each time and reach back to its current node. This movement in the clockwise direction continues until the number of agent at the current node is exactly 3, after which it changes to state Coloc.
An agent in state Gather2 (i.e., the alive agents which are not the first to detect the anomaly), first checks whether the current node has exactly 3 agents on it, if not then it moves in a counter-clockwise direction, while accompanying the pebbles present at its $home$,

### v2

@@ 9.2 Correctness and Complexity
If no agent gets destroyed by the Byzantine black hole, while performing an iteration of Algorithm 1, then our algorithm ensures that in that iteration, either the ring $R$ is explored when no agent knows the exact Byzantine black hole location or, there is one agent that knows exactly the location of the Byzantine black hole.
Proof 9.2.
Suppose till the completion of $i$-th iteration $(i>0)$ no agent is destroyed by the Byzantine black hole, then in any iteration $0<j\leq i$, the agent $a_{1}$ follows the sequence $SEQ1$, whereas $a_{2}$ follows the sequence $SEQ2$, as instructed starting from the initial $home$ node. Considering this scenario we have two cases:
  • No agents and no pebbles are destroyed: In this case, while executing the sequence $SEQ1$, the agent after changing its state to Leader from Initial, visits a new node after every 4 rounds in the clockwise direction. Since no agent has been destroyed in this iteration as well, so $a_{1}$ un-obstructively visits each node in a clockwise direction, and reaches $home$ to finally end this sequence $SEQ1$, while again changing its state to Initial. This shows that at least one agent visits each node of $R$ in an iteration completing exploration of $R$.
  • No agent gets destroyed but some pebbles are destroyed by the Byzantine black hole: Note that, this situation can only occur when the Byzantine black hole consumes the pebble carried and released by $a_{2}$, while $a_{2}$ is not at the node containing the pebble, executing the sequence $SEQ2$. It is because, in an iteration, only $a_{1}$ and $a_{2}$ moves away from $home$ while executing their respective sequences $SEQ1$ and $SEQ2$. And the pebble carried by $a_{1}$ ( i.e., the agent with the smallest ID) cannot be destroyed without destroying $a_{1}$, as $a_{1}$ always carries the pebble at each node while executing its respective sequence. So, in this situation, while $a_{2}$ ( i.e., the agent with the second lowest ID) in the state Follower-Collect moves back to the adjacent node in the counter-clockwise direction, in order to collect the already released pebble, absence of which triggers $a_{2}$ to conclude the current node to be the Byzantine black hole, and which leads the agent to immediately leave the Byzantine black hole node. In this case, the consumption of a pebble leads to the Byzantine black hole detection.
@@
If no agent gets destroyed by the Byzantine black hole while performing an iteration of Algorithm 1 but some pebbles are destroyed by the Byzantine black hole, then by Lemma 9.1, one agent (more precisely, the agent with second lowest ID, i.e., $a_{2}$) detects the exact location of the Byzantine black hole. Further, this agent continues to perpetually explore the ring $R$ by avoiding the Byzantine black hole node.
@@
    The exact location of the Byzantine black hole is detected by at least one agent.
@@
    All alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which is the Byzantine black hole.
Proof 9.4.
@@
Note that the agent that has entered the Byzantine black hole is either $a_{1}$ or $a_{2}$, based on which we have the following conditions:
@@
  • $a_{1}$ falls in to the Byzantine black hole: This situation can only occur when $a_{1}$ following sequence $SEQ1$ is in state Leader, visits the Byzantine black hole node along with its pebble. Now, based on the adversarial choice of the Byzantine black hole, the pebble may or may not be destroyed by the Byzantine black hole. Irrespective of which, $a_{2}$ within 3 additional rounds reaches the Byzantine black hole node while in the state Follower-Find, and the absence of $a_{1}$ triggers the agent $a_{2}$ to determine the current node as the Byzantine black hole, and immediately leave the current node.
@@
  • $a_{2}$ falls in to the Byzantine black hole: This can only happen when $a_{2}$ is not with $a_{1}$ at the Byzantine black hole node. Otherwise, both agents will be destroyed, contradicting our claim. Note that, $a_{2}$ can either be with the pebble or alone, while it is destroyed by the Byzantine black hole. Also note that, since $a_{2}$ is executing the sequence $SEQ2$, hence the distance between $a_{1}$ and $a_{2}$ can be at most 2. So during an iteration, if $a_{2}$ is destroyed by a Byzantine black hole, then the Byzantine black hole must be any one of the two consecutive counter-clockwise nodes, from the current position of $a_{1}$. If the current node is not $home$, in this case, $a_{1}$ gets to know about this fact, only when it finds the absence of $a_{2}$ even after waiting for 3 rounds, at the current node. Now, according to our algorithm, it leaves the pebble accompanied by it and moves from the current node in the clockwise direction to $home$ with state Report-Leader. After, reaching $home$, it waits until the end of this iteration, i.e., until $4n$ rounds from the start of this iteration and moves to state Find-Pebble. During the $14n+$ 1-th round from the start of the current iteration, $a_{3}$ finds that it is only with $a_{1}$, whereas no pebble exits, this leads the agent to change its state to Find-Pebble. In this state, both the agents move counter-clockwise from $home$ until they encounter a pebble. Note that, this pebble is the one left by $a_{1}$ after determining the fact that $a_{2}$ has entered the Byzantine black hole. So, now both these agents $a_{1}$ and $a_{3}$, declare the node with the pebble as $home$, whereas also denote the adjacent counter-clockwise node from the current node as $S_{lft}$, whereas the other node adjacent to $S_{lft}$ in the counter-clockwise direction as $S_{rgt}$. On the other hand, if the current node of $a_{1}$ is $home$, while $a_{2}$ is destroyed by the Byzantine black hole, then $a_{2}$ can never reach $home$ along with the pebble it was carrying by following the sequence $SEQ2$. So, at the end of this iteration, both $a_{1}$ and $a_{3}$ find that there are two agents (i.e., $a_{1}$ and $a_{3}$) and one pebble, respectively at $home$. This leads both of them to conclude that, the adjacent counter-clockwise node is $S_{rgt}$, whereas the adjacent counter-clockwise node of $S_{rgt}$ is $S_{lft}$.
So, in each case, we showed that either of our two conditions holds.
@@
Following from Lemma 9.3, if exactly one agent gets destroyed by the Byzantine black hole and the exact location of the Byzantine black hole is detected by at least one agent, then our algorithm ensures that the ring is perpetually explored by at least one alive agent while avoiding the Byzantine black hole node. On the contrary, all alive agents become co-located and they agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, among which one is a Byzantine black hole, then until and unless another agent falls into the Byzantine black hole, our algorithm ensures that the ring is perpetually explored, by the remaining alive agents. Note that, by Lemma 9.6 while anyone among them gets destroyed, then the other agent detects the exact Byzantine black hole node, and then further perpetually explores the ring $R$, avoiding the Byzantine black hole node.
@@
Algorithm PerpExplore-Coloc-Pbl ensures that if two among three agents get destroyed by the Byzantine black hole, then the remaining agent knows the exact location of the Byzantine black hole within some additional finite rounds without being destroyed.
Proof 9.7.
@@
Let us suppose, by contradiction, there exists a round $r$ ($>0$), within which two agents get destroyed by the Byzantine black hole, whereas the alive agent is unable to detect the exact location. Now we have the following cases:
@@
  • Both agents get destroyed by the Byzantine black hole in the same iteration: Note that in this case, the destroyed agents are $a_{1}$ and $a_{2}$, respectively, which while executing the sequences $SEQ1$ and $SEQ2$ enters the Byzantine black hole. Also, note that $a_{1}$ and $a_{2}$ can only be together at a node when $a_{2}$ has left the pebble it is carrying in the counter-clockwise node. So, now $a_{3}$ after the end of this iteration, finds that no agent among $a_{1}$ and $a_{3}$ has reached $home$, which triggers $a_{3}$ to change to state Find-BH. In this state, it moves clockwise, until it finds a pebble (this pebble is the one left by $a_{2}$ in the counter-clockwise node of the Byzantine black hole). Whenever a pebble is found, $a_{3}$ concludes that the next node is the Byzantine black hole, which contradicts our claim.
@@
  • There exists an iteration where exactly one agent is destroyed and the other one is destroyed during round $r$: Now, in this case, let $r^{\prime}<r$ be a round when exactly an agent is destroyed during an iteration, say $i>0$. Then by Lemma 9.3, within finitely many additional rounds $r_{0}$, where $r_{0}+r^{\prime}<r$, both the alive agents become co-located and agree about two consecutive nodes $S_{lft}$ and $S_{rgt}$, one among which must be the Byzantine black hole. Observe that according to our algorithm, the node that is clockwise adjacent to $S_{rgt}$ is determined to be the $home$ (if not already so). Note that in the other case according to Lemma  9.3 where after exactly one agent is destroyed by the Byzantine black hole during an iteration at a round $r^{\prime}<r$, there is an agent that knows the exact location of the Byzantine black hole within a finite additional round and the agent starts exploring the ring avoiding the Byzantine black hole and so even after round $r$ it stays alive knowing the exact location of the Byzantine black hole contradicting our assumption. Now, when both the agents are co-located during a round say $r_{0}+r^{\prime}<r$ and know that the Byzantine black hole must be any of the two consecutive nodes $S_{lft}$ or $S_{rgt}$. In this case, they move into state Detection. After which one agent moves clockwise until it reaches $S_{lft}$ and the other one moves counter-clockwise until it reaches $S_{rgt}$ and moves back to $home$. The agent that visits $S_{rgt}$ waits for the other agent at $home$ for $62n-$ 6 rounds which is sufficient for the other agent to return back at $home$. Now when during round $r$, another agent among these two gets destroyed, within at most $(2n-4)$ rounds (among the total $62n-$ 6 rounds, within which both agents are to meet) they fail to meet at $home$. Which triggers the remaining alive agent to determine exactly which one among $S_{lft}$ or $S_{rgt}$ is the Byzantine black hole. This also contradicts the claim we made earlier. Thus, the algorithm PerpExplore-Coloc-Pbl ensures that, if two among three agents get destroyed by the Byzantine black hole, the remaining agent must know the exact location of the Byzantine black hole in some finite additional rounds without getting destroyed.
10 Algorithm PerpExplore-Scat-Pbl
10.1 Detailed description and Pseudocode of the Algorithm
@@
If 4 agents are scattered in more than one nodes initially then there are 3 cases. Firstly, all four agents are at different nodes. Secondly, four agents are scattered in three different nodes nodes initially. For this case there exists exactly one node with two agents and remaining two nodes with exactly one agents each. In the third case, four agents are scattered in two nodes initially. In this case either each of the two nodes contains exactly two agents or, one node contains three agents and the other one has exactly one agent. We here describe and present the pseudocode considering the first case only. The case with 3 agents in one node can be dealt by instructing the co-located agents to execute PerpExplore-Coloc-Pbl when there are three agents on the current node (i.e., from the beginning). The algorithm we discuss here can also be used for the remaining cases with slight modifications. These modifications are described in the Remark 10.7
@@
The main idea of the PerpExplore-Scat-Pbl algorithm is that 4 agents perpetually explore the ring, when no agents are destroyed by the Byzantine black hole. Our algorithm ensures that during this exploration at most one agent can be destroyed. In this scenario, the remaining agents within further finite time, gather at a single node (which is the starting node of either of these 4 agents, hence it is safe), and after which they together start executing the algorithm PerpExplore-Coloc-Pbl.
Remark 10.1.
As mentioned in PerpExplore-Coloc-Pbl requires, 3 agents and 2 pebbles to perform PerpExploration-BBH, but in this case we instruct the agents to perform PerpExplore-Coloc-Pbl, whenever three agents and at least three pebbles are able to gather at a node. These agents performs PerpExplore-Coloc-Pbl with a slight modification that the excess pebble (or pebbles) remain at their gathered node, which they term as $home$. So, all the conditions that are mentioned, remains same in this case as well, only the condition where an agent changes to certain state $s_{i}$ by observing $t_{i}$ (say) number of pebbles at $home$, transforms to the fact that the agent changes to state $s_{i}$ by observing $t_{i}+x$ (where $x=\text{total \# pebbles}-2$) number of pebbles at $home$.
@@
At the beginning of the Algorithm 2 all agents are in state Initial1. In this state, the agents declare their current node as $home$, moreover they initialize the variables $size=0$ and $s=0$, where $size$ variable is used by the agents in order to store the length of the path between it’s own $home$ and the nearest $home$ of another agent in the clockwise direction. Further the variable $s$ indicates the state transition. If the agent is at state Forward and $s=$ 0 then the agent moved from state Initial1 to its current state (i.e., state Forward), otherwise if the agent is at state Forward and $s=1$ then this indicates it changes to state Forward from state Wait2. An agent $a_{1}$ in state Forward, initializes $W_{time}=0$ (which stores the waiting time) and checks whether $s=0$ or $s=1$. Note that, if $s=0$ and the current node is $home$, then the $a_{1}$ is first instructed to leave the pebble at current node and then move clockwise direction, after increasing $size$ by 1. The next node where $a_{1}$ has reached can either be $home$ of another agent, say $a_{2}$, or it is not $home$ for any other agents. For the first case $a_{1}$ would find a pebble on its $Current-Node$ which was left by $a_{2}$. In this case $a_{1}$ updates $s$ to 1 and moves into state Wait1. On the other hand, if the $Current-Node$ is not $home$ of any other agents then $a_{1}$ finds no pebble at the $Current-Node$ and thus moves clockwise after increasing the variable $size$ by 1. Note that, when $s=0$ and an agent is in state Forward it always finds another pebble on some node, if it is not destroyed (this pebble must be left by another agent which is simultaneously executing Algorithm 2 and currently in state Forward). This node must be the clockwise nearest $home$ (i.e., $home$ of another agent which is clockwise nearest to own $home$). Now since during each move at state Forward, an agent increases the variable $size$ by 1, the $size$ variable must store the length between an agents own $home$ and the clockwise nearest $home$ when the agent finds another pebble. Beyond this point (i.e., when $s=1$) the agents will use the $size$ variable as reference to calculate how much it should move, as after this, there is a possibility that the agent may not find a pebble at the clockwise nearest $home$.
@@
In state Wait2, the agent first checks whether $W_{time}<2n-size$, if so then it further checks whether #$\#Agent$ at current node is exactly 1. If these two conditions are satisfied then the agent just increments $W_{time}$ by 1. On the other hand, if it finds that $W_{time}<2n-size$ but #$\#Agent$ at the current node is 2, then it changes its state to Gather2 and moves clockwise and updates $W_{time}$ to 0. Otherwise, if it finds $W_{time}<2n-size$ but #$\#Agent$ at the current node is 3 then it updates $W_{time}$ to 0, whereas changes its state to Coloc. Lastly, whenever it finds $W_{time}=2n-size$, it changes its state to Forward.
[…]